# Physics-informed Value Learner for Offline Goal-Conditioned Reinforcement Learning

### v1

@@ Abstract
Offline Goal-Conditioned Reinforcement Learning (GCRL) holds great promise for domains such as autonomous navigation and locomotion, where collecting interactive data is costly and unsafe. However, it remains challenging in practice due to the need to learn from datasets with limited coverage of the state-action space and to generalize across long-horizon tasks. To improve on these challenges, we propose a Physics-informed (Pi) regularized loss for value learning, derived from the Eikonal Partial Differential Equation (PDE) and which induces a geometric inductive bias in the learned value function. Unlike generic gradient penalties that are primarily used to stabilize training, our formulation is grounded in continuous-time optimal control and encourages value functions to align with cost-to-go structures. The proposed regularizer is broadly compatible with temporal-difference-based value learning and can be integrated into existing Offline GCRL algorithms. When combined with Hierarchical Implicit Q-Learning (HIQL), the resulting method, Physics-informed HIQL (Pi-HIQL), yields significant improvements in both performance and generalization, with pronounced gains in stitching regimes and large-scale navigation tasks. Code is available at link [^1] https://github.com/VittorioGiammarino/PI-HIQL.
@@ 1 Introduction
In recent years, many of the most effective machine learning paradigms have capitalized on vast amounts of unlabeled or weakly labeled data. Similarly, in dynamic systems learning, Offline GCRL has emerged as a pivotal framework, enabling the use of large-scale, multitask datasets without requiring explicit reward annotations. Specifically, Offline RL Lange et al. (2012); Levine et al. (2020) leverages passively collected trajectories to learn control policies, offering great promise for applications such as autonomous navigation, locomotion, and manipulation, where interactive training is usually costly and unsafe. GCRL Kaelbling (1993); Schaul et al. (2015) extends this capability by enabling learning across diverse datasets without explicit rewards. Despite its potential, Offline GCRL faces significant challenges, including accurate Goal-Conditioned Value Function (GCVF) estimation from limited data, policy extraction from imperfect value functions, and generalization to unseen state-goal pairs Park et al. (2024a). Among these issues, GCVF estimation remains the most fundamental, as improvements in this area can enhance both policy extraction and generalization, ultimately advancing the entire field of Offline GCRL.
@@
Pi inductive biases offer a promising direction for improving GCVF estimation in the offline setting. As demonstrated in prior work Lien et al. (2024), Pi methods can introduce physically or geometrically meaningful structure into the learned value function, enhancing sample efficiency and generalization. In Fig. 1, we illustrate a representative GCRL task in which an agent must navigate from various starting positions in a maze to a specified goal. Fig. 1b shows the contour plot of a GCVF learned by a non-Pi state-of-the-art (SOTA) algorithm. The resulting value function fails to robustly encode obstacle constraints, leading to suboptimal policies that often fail to reach the goal. These limitations motivate the use of Pi regularizers as a principled means to incorporate structural priors into value learning, and thus improve GCVF estimation and policy reliability in complex environments.
@@
The primary contribution of this work is the introduction of a Pi Eikonal regularizer for GCVF estimation in Offline GCRL tasks. Inspired by the Eikonal PDE Noack and Clark (2017), this regularizer imposes a distance-like cost-to-go structure on the learned GCVF, serving as an effective inductive bias during training. By enforcing this structure, the regularizer improves value estimation accuracy and promotes generalization to unseen states, while also reducing the number of required training steps compared to non-Pi approaches (see Fig. 1a). In contrast to Hamilton-Jacobi-Bellman (HJB) PDE-based methods Lien et al. (2024), which require explicit system dynamics and often suffer from numerical instability Munos (1997); Munos and Bourgine (1997), our method is model-free and easy to implement. Empirically, it outperforms both HJB-regularized and unregularized baselines while adding only minimal computational overhead.
@@
To validate the effectiveness of our Eikonal regularizer, we integrate it into the HIQL framework Park et al. (2024b), a current SOTA algorithm for Offline GCRL. We refer to our method as Pi-HIQL. This choice is motivated by HIQL’s strong baseline performance, making it an ideal candidate for demonstrating the benefits of our approach. In our evaluation, conducted on the challenging OGbench benchmark Park et al. (2024c), we compare Pi-HIQL against Quasimetric RL (QRL) Wang et al. (2023), Contrastive RL (CRL) Eysenbach et al. (2022), and standard HIQL. Pi-HIQL consistently outperforms or matches the performance of HIQL and other baselines across most tasks, achieving SOTA results particularly in large-scale navigation and trajectory stitching scenarios. These gains underscore the utility of the Eikonal regularizer in enhancing GCVF estimation and overall Offline GCRL performance, with limited exceptions in tasks involving complex object interactions.
(a) Pi-HIQL (ours)(b) HIQL Park et al. (2024b)
@@
Figure 1: Contour plots of the GCVF for antmaze-giant-navigate-v0 in Park et al. (2024c), learned after 100,000 training steps by our Pi algorithm Pi-HIQL, and the standard HIQL. The plots are generated by varying the agent’s center of mass $x$-$y$ coordinates while keeping all other states fixed. Recall that the policy $\pi$ is trained to move the agent in the direction that maximizes the GCVF. The effects of the Eikonal regularizer are evident in Fig. 1a, where the contour plot closely follows the maze structure, in contrast to Fig. 1b, where the learned GCVF ignores the maze structure.
@@ 2 Related work
To the best of our knowledge, Pi regularization for value estimation has only recently been explored. Notably, Lien et al. (2024) propose an Offline RL objective derived from the HJB equation in continuous-time optimal control Munos (1997); Munos and Bourgine (1997), aiming to enforce first-order derivative consistency within the critic network. In contrast, we introduce a simpler, model-free Pi regularizer for GCVF learning, based on the residual of the Eikonal PDE. We show that this regularizer induces a distance-like structure in the value function and integrates naturally into standard GCRL pipelines.
@@
While gradient norm penalties—closely related to the Eikonal PDE residual—have been employed in generative modeling Gulrajani et al. (2017) and, more recently, in model-based RL to regularize Q-functions and mitigate overfitting Zheng et al. (2023), their application in GCRL remains, to our knowledge, unexplored. In contrast to these prior methods, which primarily aim to stabilize training, our regularizer is designed to inject a structural inductive bias into the GCVF, thereby improving sample efficiency and generalization. To the best of our knowledge, this work presents the first use of the Eikonal PDE as a regularization objective in value-based RL, and its first practical deployment in the Offline GCRL setting. More broadly, there has been a growing interest in incorporating physical priors and geometric, distance-like structures into RL algorithms, especially in model-based or Koopman-inspired frameworks Weissenbacher et al. (2022); Cheng et al. (2023). These approaches typically rely on known system dynamics or exploit symmetry groups in state space, whereas our method operates directly on the GCVF in a fully model-free setting.
@@
In parallel, the use of non-Pi constraints for value learning has been extensively studied in Offline RL. Many approaches constrain learned policies to remain close to the behavior policy, either through explicit density modeling Wu et al. (2019); Fujimoto et al. (2019); Kumar et al. (2019) or implicit divergence constraints Wang et al. (2020); Peng et al. (2019). Others directly regularize the Q-function to assign low values to out-of-distribution actions and improve robustness Kostrikov et al. (2021); Kumar et al. (2020).
@@
Our work is also closely related to the literature on GCRL Kaelbling (1993); Schaul et al. (2015). Pi-HIQL extends HIQL Park et al. (2024b) by integrating the Eikonal regularizer into the GCVF estimation loss. HIQL itself combines hierarchical actors Sutton et al. (1999a); Giammarino and Paschalidis (2021) with Implicit Q-Learning Kostrikov et al. (2021). Other GCRL methods include hindsight relabeling Andrychowicz et al. (2017), contrastive representation learning Eysenbach et al. (2022), state-occupancy matching Ma et al. (2022), and quasimetric RL Wang et al. (2023). Offline GCRL has also been studied through the lens of goal-conditioned supervised learning (GCSL) Chebotar et al. (2021); Yang et al. (2022), where goal-reaching policies are trained via conditional imitation or regression. Recent work has analyzed the out-of-distribution goal generalization problem Yang et al. (2023) and addressed GCSL via self-supervised reward shaping Mezghani et al. (2023) and occupancy-based score modeling Sikchi et al. (2023).
@@
Beyond RL, Pi losses and neural networks (NNs) have been widely applied to learn parameterizations of PDEs such as Burgers, Schrödinger, and Navier–Stokes equations Rudy et al. (2017); Raissi et al. (2019). These methods leverage automatic differentiation to estimate derivatives with respect to NN inputs and solve high-dimensional PDEs. The Eikonal PDE, in particular, has been used in seismology Smith et al. (2020) and motion planning Ni and Qureshi (2023); Ni et al. (2025), where distance fields provide essential geometric structure. In this work, we extend the use of the Eikonal PDE to the Offline GCRL domain, demonstrating how its distance-preserving properties enhance GCVF estimation, especially in large environments and when data stitching is required.
@@ 3 Preliminaries
Offline goal-conditioned reinforcement learning
@@
Hierarchical implicit Q-learning
Our algorithm, Pi-HIQL, extends the Offline GCRL algorithm HIQL, as introduced by Park et al. (2024b). HIQL incorporates two key components: a GCVF estimation process that is robust to out-of-distribution actions and a hierarchical actor. The hierarchical actor comprises a high-level policy, $\pi^{hi}_{\bm{\theta}_{hi}}:\mathcal{S}\times\mathcal{G}\to\mathcal{S}$, which predicts subgoals, and a low-level policy, $\pi^{lo}_{\bm{\theta}_{lo}}:\mathcal{S}\times\mathcal{S}\to\mathcal{P}(\mathcal{A})$, which generates actions to achieve those subgoals. Robustness in the GCVF estimation step is enabled by an action-free variant of implicit $Q$-learning Kostrikov et al. (2021), inspired by Xu et al. (2022); Ghosh et al. (2023):
@@
where $\bar{\bm{\theta}}_{V}$ denotes the parameters of the target GCVF network Mnih (2013) and $L_{2}^{\iota}(\cdot)$ represents the expectile loss function with $\iota\in[0.5,1]$: $L_{2}^{\iota}(x)=|\iota-\mathbbm{1}(x<0)|x^{2}$. In Eq. (1), the expectile regression induced by $L_{2}^{\iota}(\cdot)$ replaces the $\max$ operator in the Bellman equation Sutton and Barto (2018) with the goal of avoiding queries of out-of-distributions actions. The ability to properly handle overestimated values for out-of-distribution actions is a critical challenge in Offline RL, since, unlike in online RL, erroneous estimates cannot be corrected through environment interactions. The estimated GCVF is subsequently used to train the hierarchical actor where $\pi^{hi}_{\bm{\theta}_{hi}}(s_{t+k}|s_{t},g)$ and $\pi^{lo}_{\bm{\theta}_{lo}}(a|s_{t},s_{t+k})$ aim to maximize $V_{\bm{\theta}_{V}}(s_{t+k},g)$ and $V_{\bm{\theta}_{V}}(s_{t+1},s_{t+k})$, respectively. It is shown in Park et al. (2024b) that this hierarchy, compared to the flat formulation using a single policy $\pi_{\bm{\theta}}(a|s_{t},g)$, can better address low signal-to-noise ratios in the estimated GCVF.
@@
The Eikonal equation is a non-linear first-order PDE that describes wave propagation in heterogeneous media Noack and Clark (2017). It is expressed in its general form as:
@@
where $||\cdot||$ denotes the Euclidean norm, $T:\mathcal{S}\times\mathcal{G}\to\mathbb{R}$ represents the travel-time through the medium from the state $s$ to a goal location $g$ and $\nabla_{s}T(s,g)$ is the partial derivative of the travel-time $T$ with respect to $s$. The function $S:\mathcal{S}\to\mathbb{R}$ defines the speed profile of the medium in the state location $s$. As described in Ni and Qureshi (2023), higher values for $S(s)$ lead to a low travel-time $T(s,g)$ from $s$ to $g$ and therefore to preferable paths $(s_{0},s_{1}\dots,s_{T})$ compared to those with lower $S(s)$. Consequently, the solution to the Eikonal PDE in (2), represented by the travel-time $T(s,g)$, reflects a cost-to-go structure. Minimizing $T(s,g)$ yields the shortest travel time from $s$ to $g$, as determined by the speed profile $S(s)$. In other words, $T(s,g)$ encodes a GCVF for a specific class of optimal control problems. We formally establish this connection in the next section. In our method, we propose a Pi regularizer, inspired by the Eikonal PDE in (2), with the goal of providing an additional distance-like structure to the learned GCVF. Under smooth dynamics assumptions, we demonstrate that our regularizer significantly improves performance over the SOTA baselines while adding minimal complexity.
@@ 4 Physics-informed Eikonal regularizer
In this section, we first relate the Eikonal PDE in (2) to the HJB equation Munos (1997); Munos and Bourgine (1997) for continuous-time, undiscounted optimal control. This connection draws parallels to prior studies such as Lien et al. (2024), offering additional motivation for our regularizer and insights into its empirical effectiveness. We then introduce the Eikonal-regularized loss for GCVF learning, which, when combined with a hierarchical actor, forms the core of our Pi-HIQL algorithm. Finally, we summarize the main components of Pi-HIQL.
Optimal control perspective on the Eikonal equation
We start our analysis considering the following continuous-time dynamical system:
@@
where $s(t)$ denotes the state of the system at time $t$, $\dot{s}(t)$ its derivative and $a(t)$ the control action. The function $f(\cdot)$ represents the system dynamics, determining how the state evolves in response to the control action. As common in the literature, we assume $f(\cdot)$ to be a Lipschitz continuous function Kim et al. (2021). Given the initial conditions $s(0)=s_{0}$, $a(0)=a_{0}$ and the goal $s(T)=g$, the undiscounted optimal control problem seeks to minimize
@@
[…]
@@
where $\Delta t$ is a small time step. Note that, unless otherwise specified, throughout this section we keep the standard continuous-time optimal control theory notation, where $V(s,g)$ is used in place of $V^{*}(s,g)$ to denote the optimal value function Fleming and Soner (2006).
@@
which holds true at optimality. The HJB equation in Eq. 5 encodes the relationship between the system dynamics, the cost function, and the value function; where the left-hand side $H(s,g,\nabla_{s}V(s,g))\equiv\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\intercal}f(s,a)]$ is referred to as the Hamiltonian. The following proposition establishes the connection between the Hamiltonian and the Eikonal PDE in (2), highlighting their equivalence under specific conditions.
@@
Temporal-difference learning (e.g., (1)) is known to converge to the optimal GCVF $V^{*}$ under ideal conditions—namely, on-policy data and an infinite sample budget Sutton and Barto (2018). However, these assumptions are often violated in the offline setting, where biased datasets and long-horizon tasks exacerbate extrapolation error and limit generalization Park et al. (2024b). In this context, the Eikonal residual in Eq. (7) can play a crucial role by introducing a geometric inductive bias that encourages the learned value function to behave like a distance field, through the constraint $\|\nabla_{s}V_{\bm{\theta}}(s,g)\|\approx c(s)$. This regularization is particularly effective when the true $V^{*}$ is Lipschitz continuous—i.e., in environments where the dynamics do not induce sharp discontinuities Banjanin and Burden (2019). Under such conditions, the Eikonal residual shapes the gradient norm of the learned GCVF $V_{\bm{\theta}}$ to match the local structure of $V^{*}$, up to a scaling factor. This effect is supported by Rademacher’s theorem Heinonen (2005), which guarantees that Lipschitz continuous functions are differentiable almost everywhere, with $\|\nabla_{s}V^{*}(s,g)\|\leq L$, where $L$ is the minimal Lipschitz constant. As a result, when $V^{*}$ is Lipschitz continuous, the Eikonal residual in Eq. (7) provides a principled inductive bias that, as we empirically demonstrate, improves both sample efficiency and generalization—particularly in offline, long-horizon GCRL tasks with smooth dynamics.
Furthermore, in practice, the Eikonal residual offers a tractable alternative to HJB regularization in model-free settings where the dynamics $f(s,a)$ are unknown. Prior work Lien et al. (2024) approximates the HJB term using finite differences, i.e., $f(s,a)\approx s^{\prime}-s$, but our experiments show no clear advantage over the simpler Eikonal residual regularization. Moreover, note that, while the HJB equation in Eq. (5) holds only at optimality, the inequality in Proposition 4.1 remains valid throughout training, providing consistent structural guidance even when $V$ is still suboptimal.
@@
where the term in (8) corresponds to the expectile regression in (1) (cf. Park et al. (2024b)) and (9) is our Eikonal regularizer. In (9), $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)||$ is the Euclidean norm of the GCVF gradient with respect to its input $s$ and $S(s)$ is a pre-specified speed profile that maps states to scalar values (see Preliminaries in Section 3). The term $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)$ is computed via automatic differentiation, following standard approaches in PiNNs (see Algorithm 2 in Appendix C). The speed profile $S(s)$ is designed such that it encapsulates additional biases or task-specific information into the Eikonal regularizer Ni and Qureshi (2023). We demonstrate in our experiments that the simple choice of $S(s)=1$ works best in practice; however, we believe that more interesting designs might further improve collision avoidance in cluttered environments and consequently enhance both safety and performance. We defer this direction on Pi Safe GCRL to future work. Finally, note that in (9), with respect to the upper bound in (6), we set $c^{*}(s)=-1$ and opt to multiply $S(s)$ with $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)||$ rather than using $c^{*}(s)=-1/S(s)$ as in (2). This in order to ensure numerical stability.

### v3

@@ Abstract
Offline Goal-Conditioned Reinforcement Learning (GCRL) holds great promise for domains such as autonomous navigation and locomotion, where collecting interactive data is costly and unsafe. However, it remains challenging in practice due to the need to learn from datasets with limited coverage of the state-action space and to generalize across long-horizon tasks. To improve on these challenges, we propose a Physics-informed (Pi) regularized loss for value learning, derived from the Eikonal Partial Differential Equation (PDE) and which induces a geometric inductive bias in the learned value function. Unlike generic gradient penalties that are primarily used to stabilize training, our formulation is grounded in continuous-time optimal control and encourages value functions to align with cost-to-go structures. The proposed regularizer is broadly compatible with temporal-difference-based value learning and can be integrated into existing Offline GCRL algorithms. When combined with Hierarchical Implicit Q-Learning (HIQL), the resulting method, Eikonal-regularized HIQL (Eik-HIQL), yields significant improvements in both performance and generalization, with pronounced gains in stitching regimes and large-scale navigation tasks. Code is available at link [^1] https://github.com/VittorioGiammarino/Eik-HIQL.
@@ 1 Introduction
In recent years, many of the most effective machine learning paradigms have capitalized on vast amounts of unlabeled or weakly labeled data. Similarly, in dynamic systems learning, Offline Goal-Conditioned Reinforcement Learning (GCRL) has emerged as a pivotal framework, enabling the use of large-scale, multitask datasets without requiring explicit reward annotations. Specifically, Offline RL [1, 2] leverages passively collected trajectories to learn control policies, offering great promise for applications such as autonomous navigation, locomotion, and manipulation, where interactive training is usually costly and unsafe. GCRL [3, 4] extends this capability by enabling learning across diverse datasets without explicit rewards. Despite its potential, Offline GCRL faces significant challenges, including accurate Goal-Conditioned Value Function (GCVF) estimation from limited data, policy extraction from imperfect value functions, and generalization to unseen state-goal pairs [5]. Among these issues, GCVF estimation remains the most fundamental, as improvements in this area can enhance both policy extraction and generalization, ultimately advancing the entire field of GCRL.
@@
Physics-informed (Pi) inductive biases, defined as structural constraints grounded in physical laws such as symmetry, conservation principles, or consistency with Partial Differential Equations (PDEs), provide a promising direction for enhancing GCVF estimation in the offline setting. As demonstrated in prior work [6], Pi methods can introduce physically or geometrically meaningful structure into the learned value function, enhancing both sample efficiency and generalization. In Fig. 1, we illustrate a representative GCRL task in which an agent must navigate from various starting positions in a maze to a specified goal. Fig. 1b shows the contour plot of a GCVF learned by a non-Pi state-of-the-art (SOTA) algorithm. The resulting value function fails to robustly encode obstacle constraints, leading to suboptimal policies that often fail to reach the goal. These limitations motivate the use of Pi regularizers as a principled means to incorporate structural priors into value learning, and thus improve GCVF estimation in complex environments.
@@
The primary contribution of this work is the introduction of an Eikonal regularizer for GCVF estimation in Offline GCRL tasks. Inspired by the Eikonal PDE [7], this regularizer imposes a distance-like cost-to-go structure on the learned GCVF, serving as an effective inductive bias during training. By enforcing this structure, the regularizer improves value estimation accuracy and promotes generalization to unseen states, while also reducing the number of required training steps compared to non-Pi approaches (see Fig. 1a). In contrast to Hamilton-Jacobi-Bellman (HJB) PDE-based methods [6], which require explicit system dynamics and often suffer from numerical instability [8, 9], our method is model-free and easy to implement. Empirically, it outperforms both HJB-regularized and unregularized baselines while adding only minimal computational overhead.
@@
To validate the effectiveness of our Eikonal regularizer, we integrate it into the Hierarchical Implicit Q-Learning (HIQL) framework [10], a SOTA algorithm for Offline GCRL. We refer to this variant as Eik-HIQL. This choice is motivated by HIQL’s strong baseline performance, making it an ideal candidate to highlight the benefits of our approach. Importantly, the Eikonal regularizer is broadly compatible with other temporal-difference-based algorithms, as we further demonstrate empirically in Appendix F. Our evaluation, conducted on the challenging OGbench benchmark [11], compares Eik-HIQL against Quasimetric RL (QRL) [12], Contrastive RL (CRL) [13], and the standard HIQL baseline. Eik-HIQL consistently outperforms or matches the baselines, achieving SOTA results in large-scale navigation and trajectory stitching scenarios. These gains underscore the utility of the Eikonal regularizer in enhancing GCVF estimation and overall Offline GCRL performance, with limited exceptions in tasks involving complex object interactions.
(a) Eik-HIQL (ours)(b) HIQL [10]
@@
Figure 1: Contour plots of the GCVF for antmaze-giant-navigate-v0 in [11], learned after 100,000 training steps by our Physics-informed algorithm Eik-HIQL, and the standard HIQL. The plots are generated by varying the agent’s center of mass $x$-$y$ coordinates while keeping all other states fixed. Recall that the policy $\pi$ is trained to move the agent in the direction that maximizes the GCVF. The effects of the Eikonal regularizer are evident in Fig. 1a, where the contour plot closely follows the maze structure, in contrast to Fig. 1b, where the learned GCVF ignores the maze structure.
@@ 2 Related work
To the best of our knowledge, Pi regularization for value estimation has only recently been explored. Notably, Lien et al. [6] propose an Offline RL objective derived from the HJB equation in continuous-time optimal control [8, 9], aiming to enforce first-order derivative consistency within the critic network. In contrast, we introduce a simpler, model-free Pi regularizer for GCVF learning, based on the residual of the Eikonal PDE. We show that this regularizer induces a distance-like structure in the value function and integrates naturally into standard temporal-difference-based GCRL pipelines.
@@
While gradient norm penalties, closely related to the Eikonal PDE residual, have been employed in generative modeling [14] and, more recently, in model-based RL to regularize Q-functions and mitigate overfitting [15], their application in GCRL remains, to our knowledge, unexplored. In contrast to these prior methods, which primarily aim to stabilize training, our regularizer is designed to inject a structural inductive bias into the GCVF, thereby improving sample efficiency and generalization. To the best of our knowledge, this work presents the first use of the Eikonal PDE as a regularization objective in value-based RL, and its first practical deployment in the Offline GCRL setting. More broadly, there has been a growing interest in incorporating physical priors and geometric, distance-like structures into RL algorithms, especially in model-based or Koopman-inspired frameworks [16, 17]. These approaches typically leverage the Koopman operator, which assumes access to (or approximations of) the underlying system dynamics, and may incorporate structural information such as reversibility or symmetry of the dynamics. In contrast, our method operates directly on the GCVF in a fully model-free setting.
@@
In parallel, the use of non-Pi constraints for value learning has been extensively studied in Offline RL. Many approaches constrain learned policies to remain close to the behavior policy, either through explicit density modeling [18, 19, 20] or implicit divergence constraints [21, 22]. Others directly regularize the Q-function to assign low values to out-of-distribution actions and improve robustness [23, 24].
@@
Our work is also closely related to the literature on GCRL [3, 4]. Eik-HIQL extends HIQL [10] by integrating the Eikonal regularizer into the GCVF estimation loss. HIQL itself combines hierarchical actors [25, 26] with Implicit Q-Learning [23]. Other GCRL methods include hindsight relabeling [27], contrastive representation learning [13], state-occupancy matching [28], and quasimetric RL [12]. Offline GCRL has also been studied through the lens of goal-conditioned supervised learning (GCSL) [29, 30], where goal-reaching policies are trained via conditional imitation or regression. Recent work has analyzed the out-of-distribution goal generalization problem [31] and addressed GCSL via self-supervised reward shaping [32] and occupancy-based score modeling [33].
@@
Beyond RL, Pi losses and neural networks (NNs) have been widely applied to learn parameterizations of PDEs such as Burgers, Schrödinger, and Navier–Stokes equations [34, 35]. These methods leverage automatic differentiation to estimate derivatives with respect to NN inputs and solve high-dimensional PDEs. The Eikonal PDE, in particular, has been used in seismology [36] and motion planning [37, 38], where distance fields provide essential geometric structure. In this work, we extend the use of the Eikonal PDE to the Offline GCRL domain, demonstrating how its distance-preserving properties enhance GCVF estimation, especially in large environments and when data stitching is required.
@@ 3 Preliminaries
Offline GCRL
@@
Hierarchical IQL
Our algorithm, Eik-HIQL, extends the Offline GCRL algorithm HIQL, as introduced by Park et al. [10]. HIQL incorporates two key components: a GCVF estimation process that is robust to out-of-distribution actions and a hierarchical actor. The hierarchical actor comprises a high-level policy, $\pi^{hi}_{\bm{\theta}_{hi}}:\mathcal{S}\times\mathcal{G}\to\mathcal{S}$, which predicts subgoals, and a low-level policy, $\pi^{lo}_{\bm{\theta}_{lo}}:\mathcal{S}\times\mathcal{S}\to\mathcal{P}(\mathcal{A})$, which generates actions to achieve those subgoals. Robustness in the GCVF estimation step is enabled by an action-free variant of implicit $Q$-learning [23], inspired by [39, 40]:
@@
where $\bar{\bm{\theta}}_{V}$ denotes the parameters of the target GCVF network [41] and $L_{2}^{\iota}(\cdot)$ represents the expectile loss function with $\iota\in[0.5,1]$: $L_{2}^{\iota}(x)=|\iota-\mathbbm{1}(x<0)|x^{2}$. In Eq. (1), the expectile regression induced by $L_{2}^{\iota}(\cdot)$ replaces the $\max$ operator in the Bellman equation [42] with the goal of avoiding queries of out-of-distributions actions. The ability to properly handle overestimated values for out-of-distribution actions is a critical challenge in Offline RL, since, unlike in online RL, erroneous estimates cannot be corrected through environment interactions. The estimated GCVF is subsequently used to train the hierarchical actor where $\pi^{hi}_{\bm{\theta}_{hi}}(s_{t+k}|s_{t},g)$ and $\pi^{lo}_{\bm{\theta}_{lo}}(a|s_{t},s_{t+k})$ aim to maximize $V_{\bm{\theta}_{V}}(s_{t+k},g)$ and $V_{\bm{\theta}_{V}}(s_{t+1},s_{t+k})$, respectively. It is shown in Park et al. [10] that this hierarchy, compared to the flat formulation using a single policy $\pi_{\bm{\theta}}(a|s_{t},g)$, can better address low signal-to-noise ratios in the estimated GCVF.
@@
The Eikonal equation is a non-linear first-order PDE that describes wave propagation in heterogeneous media [7]. It is expressed in its general form as:
@@
where $||\cdot||$ denotes the Euclidean norm, $T:\mathcal{S}\times\mathcal{G}\to\mathbb{R}$ represents the travel-time through the medium from the state $s$ to a goal location $g$ and $\nabla_{s}T(s,g)$ is the partial derivative of the travel-time $T$ with respect to $s$. The function $S:\mathcal{S}\to\mathbb{R}$ defines the speed profile of the medium in the state location $s$. As described in [37], higher values for $S(s)$ lead to a low travel-time $T(s,g)$ from $s$ to $g$ and therefore to preferable paths $(s_{0},s_{1}\dots,s_{T})$ compared to those with lower $S(s)$. Consequently, the solution to the Eikonal PDE in (2), represented by the travel-time $T(s,g)$, reflects a cost-to-go structure. Minimizing $T(s,g)$ yields the shortest travel time from $s$ to $g$, as determined by the speed profile $S(s)$. In other words, $T(s,g)$ encodes a GCVF for a specific class of optimal control problems. We formally establish this connection in the next section. In our method, we propose a regularizer, inspired by the Eikonal PDE in (2), with the goal of providing an additional distance-like structure to the learned GCVF. Under smooth dynamics assumptions, we demonstrate that our regularizer significantly improves performance over the SOTA baselines while adding minimal complexity.
@@ 4 Physics-informed Eikonal regularizer
In this section, we first relate the Eikonal PDE in (2) to the HJB equation [8, 9] for continuous-time, undiscounted optimal control. This connection draws parallels to prior studies such as Lien et al. [6], offering additional motivation for our regularizer and insights into its empirical effectiveness. We then introduce the Eikonal-regularized loss for GCVF learning, which, when combined with a hierarchical actor, forms the core of our Eik-HIQL algorithm. Finally, we summarize the main components of Eik-HIQL.
Optimal control perspective on the Eikonal PDE
We start our analysis by considering the following continuous-time dynamical system:
@@
where $s(t)$ denotes the state of the system at time $t$, $\dot{s}(t)$ its derivative and $a(t)$ the control action. The function $f(\cdot)$ represents the system dynamics, determining how the state evolves in response to the control action. As common in the literature, we assume $f(\cdot)$ to be a Lipschitz continuous function [43]. Given the initial conditions $s(0)=s_{0}$, $a(0)=a_{0}$ and the goal $s(T)=g$, the undiscounted optimal control problem seeks to minimize
@@
where $c(s(t),a(t))$ is the instantaneous cost function. The optimal value function $V(s,g)$ associated with (3) is defined as
@@
where $\Delta t$ is a small time step. Note that, unless otherwise specified, throughout this section we keep the standard continuous-time optimal control theory notation, where $V(s,g)$ is used in place of $V^{*}(s,g)$ to denote the optimal value function [44].
@@
which holds true at optimality. Refer to Appendix C for the step-by-step derivations. The HJB equation in (5) encodes the relationship between the system dynamics, the cost function, and the value function; where the left-hand side $H(s,g,\nabla_{s}V(s,g))\equiv\inf_{a\in\mathcal{A}}[c(s,a)+\nabla_{s}V(s,g)^{\intercal}f(s,a)]$ is referred to as the Hamiltonian. The following proposition establishes the connection between the Hamiltonian and the Eikonal PDE in (2), highlighting their equivalence under specific conditions.
@@
Temporal-difference learning (e.g., (1)) is known to converge to the optimal GCVF $V^{*}$ under ideal conditions: namely, on-policy data and an infinite sample budget [42]. However, these assumptions are often violated in the offline setting, where biased datasets and long-horizon tasks exacerbate extrapolation error and limit generalization [10]. In this context, the Eikonal residual in Eq. (7) can play a crucial role by introducing a geometric inductive bias that encourages the learned value function to behave like a distance field, through the constraint $\|\nabla_{s}V_{\bm{\theta}}(s,g)\|\approx c(s)$. This regularization is particularly effective when the true $V^{*}$ is Lipschitz continuous, i.e., in environments where the dynamics do not induce sharp discontinuities [45]. Under such conditions, the Eikonal residual shapes the gradient norm of the learned GCVF $V_{\bm{\theta}}$ to match the local structure of $V^{*}$, up to a scaling factor. This effect is supported by Rademacher’s theorem [46], which guarantees that Lipschitz continuous functions are differentiable almost everywhere, with $\|\nabla_{s}V^{*}(s,g)\|\leq L$, where $L$ is the minimal Lipschitz constant. As a result, when $V^{*}$ is Lipschitz continuous, the Eikonal residual in Eq. (7) provides a principled inductive bias that, as we empirically demonstrate, improves both sample efficiency and generalization, particularly in long-horizon tasks with smooth dynamics.
Furthermore, in practice, the Eikonal residual offers a tractable alternative to HJB regularization in model-free settings where the dynamics $f(s,a)$ are unknown. Prior work [6] approximates the HJB term using finite differences, i.e., $f(s,a)\approx s^{\prime}-s$, but our experiments show no clear advantage over the simpler Eikonal residual regularization. Moreover, note that, while the HJB equation in Eq. (5) holds only at optimality, the inequality in Proposition 4.1 remains valid throughout training, providing consistent structural guidance even when $V$ is suboptimal.
@@
where the term in (8) corresponds to the expectile regression in (1) (cf. [10]) and (9) is our Eikonal regularizer. In (9), $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)||$ is the Euclidean norm of the GCVF gradient with respect to its input $s$ and $S(s)$ is a pre-specified speed profile that maps states to scalar values (see Preliminaries in Section 3). The term $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)$ is computed via automatic differentiation, following standard approaches in PiNNs (see Algorithm 2 in Appendix D). The speed profile $S(s)$ is designed such that it encapsulates additional biases or task-specific information into the Eikonal regularizer [37]. We demonstrate in our experiments that the simple choice of $S(s)=1$ works best in practice; however, we believe that more interesting designs might further improve collision avoidance in cluttered environments and consequently enhance both safety and performance. We defer this direction on Pi Safe GCRL to future work. Finally, note that in (9), with respect to the upper bound in (6), we set $c^{*}(s)=-1$ and opt to multiply $S(s)$ with $\nabla_{s}V_{\bm{\theta}_{V}}(s,g)||$ rather than using $c^{*}(s)=-1/S(s)$ as in (2). This in order to ensure numerical stability.